# Noncoincidence $f(Q)$-Cosmology with Dark Matter Coupled to Gravity

### v1

@@ I Introduction
In modern cosmology, where spacetime is usually modeled by the isotropic and homogeneous Friedmann-Lemaître-Robertson-Walker (FLRW) geometry, there exist four possible connections consistent with being symmetric, flat, and compatible with the isometries of the background Hohmann ; fq4 . Three of these correspond to the spatially flat case, while the fourth applies to the nonzero-curvature case. The associated degrees of freedom of $f\left(Q\right)$ theory have been investigated in detail in the literature. In particular, the field equations admit a minisuperspace description an3 . It was found that when the connection is defined in the noncoincidence gauge, additional dynamical degrees of freedom appear, which can be interpreted as scalar fields an2 . On the other hand, when the coincidence gauge is adopted, the resulting cosmological equations coincide with those of $f\left(T\right)$ teleparallel gravity tel1 .
@@
In ppr1 ; ppr2 it was shown that $f\left(Q\right)$ gravity suffers from strong coupling or the appearance of ghosts at the level of cosmological perturbations. Nevertheless, these problems can be avoided when the matter sector is coupled to the gravitational field directly in the Action Integral. This type of interaction is the focus of the present work, where we restrict attention to the background dynamics. Specifically, we perform a detailed phase-space analysis of the cosmological field equations in the presence of a coupling function. The stationary points are identified and used to reconstruct the possible cosmological histories supported by the theory.
@@ III Matter Coupled to Gravity
such that in the limit $f(Q)\sim Q$, general relativity is recovered. Such interacting models have been examined before in other modified theories of gravity, see for instance fr1 ; fr2 ; fr3 . In our consideration the interacting function has been introduced in a way such that to introduce the minimum nonlinear terms in the cosmological field equations.
@@ IV Exact Solutions
We examine the existence of two exact cosmological solutions of special interests, the power-law solution, $a\left(t\right)=a_{0}t^{p}$, with Hubble function $H\left(t\right)=\frac{p}{t}$, and the de Sitter solution $a\left(t\right)=a_{0}e^{H_{0}t}$ with constant Hubble function $H\left(t\right)=H_{0}$. The power-law solution can describe a radiation epoch, for $p=\frac{1}{2}$, or the matter dominated era, for $p=\frac{2}{3}$.
@@ VI Conclusions
We investigate the phase space of the resulting dynamical system and we explore the existence of stationary points. The stationary points identified in the phase-space analysis represent distinct cosmological regimes, each characterized by specific scalar field dynamics and energy contributions. However, their physical relevance depends critically on their stability properties. For a generic function $f\left(Q\right)$, we determined three families of stationary points, which describe de Sitter solutions where the potential term dominates, points AA and $D$; a matter-dominated epoch described by point $B$, and scaling solutions given by the stationary points $C$ and $E$.

### v2

@@ I Introduction
In modern cosmology, where spacetime is usually modeled by the isotropic and homogeneous Friedmann-Lemaître-Robertson-Walker (FLRW) geometry, there exist four possible connections consistent with being symmetric, flat, and compatible with the isometries of the background Hohmann ; fq4 . Three of these correspond to the spatially flat case, while the fourth applies to the nonzero-curvature case. The associated degrees of freedom of $f\left(Q\right)$ theory have been investigated in detail in the literature. In particular, the field equations admit a minisuperspace description an3 . It was found that when the connection is defined in the noncoincidence gauge, additional dynamical degrees of freedom appear, which can be interpreted as scalar fields an2 . On the other hand, when the coincidence gauge is adopted, the resulting cosmological equations coincide with those of $f\left(T\right)$ teleparallel gravity tel1 . Although torsion teleparallel theory of gravities, can violate the local Lorentz symmetry in some extensions, for details we refer to the discussion in lo1 ; lo2 ; lo3 , this is not the case of symmetric teleparallel theory.
@@
In ppr1 ; ppr2 it was shown that $f\left(Q\right)$ gravity suffers from strong coupling or the appearance of ghosts at the level of cosmological perturbations. Nevertheless, these problems can be avoided when the matter sector is coupled to the gravitational field directly in the Action Integral. This type of interaction is the focus of the present work, where we restrict attention to the background dynamics. Specifically, we perform a detailed phase-space analysis of the cosmological field equations in the presence of a coupling function. The stationary points are identified and used to reconstruct the possible cosmological histories supported by the theory. Cosmological models with an interacting dark sector have drawn the attention recently because they can describe the recent cosmological data dt1 ; dt2 ; dt3 ; dt4 . It was found that the late-time universe indicate a strong interaction coupling between the dark energy and the dark matter dt5 ; dt6 ; dt7 . In $f\left(Q\right)$ gravity, dark energy has a geometric origin, thus in order to introduce the interaction with the dark matter component, we consider that the Lagrangian function of the latter is coupled with a nonlinear function to the nonmetricity scalar $Q$.
@@ III Matter Coupled to Gravity
such that in the limit $f(Q)\sim Q$, general relativity is recovered. Such interacting models have been examined before in other modified theories of gravity, see for instance fr1 ; fr2 ; fr3 . In our consideration the interacting function has been introduced in a way such that to introduce the minimum nonlinear terms in the cosmological field equations. The form of the coupling function has been considered such that to have the minimum number of degrees of freedom introduced by the coupling function and have similarities with other theories of gravity such as the Weyl Integrable Spacetime or the Chameleon Mechanism.
@@ IV Exact Solutions
We examine the existence of two exact cosmological solutions of special interests, the power-law solution, $a\left(t\right)=a_{0}t^{p}$, with Hubble function $H\left(t\right)=\frac{p}{t}$, and the de Sitter solution $a\left(t\right)=a_{0}e^{H_{0}t}$ with constant Hubble function $H\left(t\right)=H_{0}$. The power-law solution can describe a radiation epoch, for $p=\frac{1}{2}$, or the matter dominated era, for $p=\frac{2}{3}$. This analysis will reveal important information about the behaviour of the $f\left(Q\right)$-function regarding the existence of these main eras.
@@ VI Conclusions
We investigate the phase space of the resulting dynamical system and we explore the existence of stationary points. The stationary points identified in the phase-space analysis represent distinct cosmological regimes, each characterized by specific scalar field dynamics and energy contributions. However, their physical relevance depends critically on their stability properties. For a generic function $f\left(Q\right)$, we determined three families of stationary points, which describe de Sitter solutions where the potential term dominates, points $A$ and $D$; a matter-dominated epoch described by point $B$, and scaling solutions given by the stationary points $C$ and $E$.
@@ Acknowledgements.
Data Availability Statement: Data sharing not applicable- no new data generated, the article describes entirely theoretical research.